# Sea-Land-Cloud Segmentation in Satellite Hyperspectral Imagery by Deep Learning

### v2

@@ 2.3.1 Processing of 1D vs 3D Data
The ML and DL models presented in this work handle the data in either one-dimensional (1D) or three-dimensional (3D) formats. On the one hand, when we refer to 1D processing, it implies that the models make predictions solely based on spectral characteristics, disregarding any nearby spatial context. Examples of such models include 1D-CNNs based on 1D convolutions, where the input consists of a spectral signature with no spatial context. Henceforth, we refer to these types of models as 1D models. However, when we refer to 3D processing, it signifies that the models concurrently analyze multiple adjacent spatial pixels, incorporating the hyperspectral datacube with both spectral (1D) and spatial (2D) information to predict the class for each pixel. Models like 2D-CNNs, that we refer henceforth as 2D models serve as examples, where the input is a 3D image containing both spatial and spectral data. Convolutional layers in 2D-CNN models operate by first applying 2D convolutions individually between each spectral channel and its respective neural kernel component. Subsequently, they aggregate the 2D convolutional results stacked in the spectral dimension to produce a final 2D feature map. Finally, we describe in B the arrangement of the data dimensions required for training and inference. In the appendix, we offer e.g. information about how the images are divided into 3D patches for 2D-CNNs and the size of the patches used.
2.3.2 Baseline Segmentation Models
@@
As a starting point, we begin using simple baseline models consisting of traditional ML algorithms as they are often more interpretable than deep neural networks. For convenience, during this initial phase, the models process 1D data (spectral signatures), which is a simpler processing strategy compared to handling 3D image patches. The initial ML baseline consists of Stochastic Gradient Descent (referred henceforth as 1D-ML/SGD), Gaussian Naïve Bayes (1D-ML/NB), Linear Discriminant Analysis (1D-ML/LDA), and Quadratic Discriminant Analysis (1D-ML/QDA). Next, we incorporate multiple state-of-the-art DL models for 3D patch processing, such as NU-Net [38] and NU-Net-mod [13], previously employed in satellite imagery for cloud detection in RGB (FACSAT-1) and MSI (Sentinel-2B). Furthermore, we incorporate additional DL models, namely C-UNet and C-UNet++, distinguished by the authors [39] as lightweight for their minimal parameter count and low computational requirements, making them well-suited for deployment on satellites. C-UNet and C-UNet++ are compact versions of the standard encoder-decoder U-Net architecture [40]. The subsequent literature [41] building upon [39], explores C-UNet and C-UNet++ models suggesting their possible use for segmentation in HS missions such as HYPSO-1. However, the method in [41] involves training on only two limited HS images, the conventional Pavia Centre and AeroRIT scene, which use sensors different from HYPSO-1 and do not include any cloud or oceanic data, which is of particular interest for ocean observation in HYPSO-1. In our work, we refer to the compact models in [41] as 2D-CUNet and 2D-CUNet++, and to their further reduced versions in [41] as 2D-CUNet Reduced and 2D-CUNet++ Reduced. Moreover, we add models from [42], which [42] refers to as the FAUBAI model with reference to the FAUBAI project that focuses on segmentation for HS imagery for its possible future use in satellites. We refer to the networks in [42] as 2D-UNet FAUBAI and to its reduced version as 2D-UNet FAUBAI Reduced. Although Netteland [42] recommends any of these two networks for tasks such as cloud segmentation in HS satellite imagery, we note that similarly to [41], Netteland [42] conducted model testing only on a restricted set of four conventional HS images, namely, Pavia Centre, Pavia University, Salinas, and Indian Pines, which do not contain cloud or oceanic data.
2.3.3 Customised Segmentation Models
2.3.3.1 Review of 1D-CNNs: Liuetal, Huetal, and LucasCNN
@@
The Liuetal [43], Huetal [44] and LucasCNN [45] 1D-CNN models have been used before for data analysis in soil spectroscopy to segment various soil texture classes in HS data. Additionally, in our previous article [34], we verified that a modified version of Liuetal achieved high performance in satellite HS imagery, surpassing the state-of-the-art in soil spectroscopy [45]. In this work, we find Liuetal, Huetal, and LucasCNN of interest also for in-orbit inference due to their simple 1D processing approach. Therefore, assuming 112 spectral channels and 3 classes to segment, in Table I we show the parameter counts for the three models. The parameter counts are in the order of tens of thousands with the Liuetal having the lowest number of parameters among the three. Following our pursuit of further reducing the parameter count to merely a few thousands, more convenient for inference aboard satellites, in the next sections we accordingly modify the soil models in [45], while aiming to preserve their performance.
@@
Liuetal 22,755 Hidden: 4 1D Convolutions & 4 Max Poolings Output: Flatten & 1 Dense Hidden/Convolutions: N=32,32,64,64; K=3; Act.=ReLU Hidden/Poolings: ÷2 Output/Dense: N=3; Act.=Softmax
1D-Justo-LiuNet 4,563 Hidden: 4 1D Convolutions & 4 Max Poolings Output: Flatten & 1 Dense Hidden/Convolutions: N=6,12,18,24; K=6; Act.=ReLU Hidden/Poolings: ÷2 Output/Dense: N=3; Act.=Softmax
Huetal 100,663 Hidden: 1 1D Convolution & 1 Max Pooling Output: Flatten & 2 Dense Hidden/Convolution: N=20; K=12; Act.=Tanh Hidden/Pooling: ÷2 Output/Dense 1: N=100; Act.=Tanh Output/Dense 2: N=3; Act.=Softmax
1D-Justo-HuNet 9,543 Hidden: 1 1D Convolution & 1 Max Pooling Output: Flatten & 2 Dense Hidden/Convolution: N=6; K=9; Act.=Tanh Hidden/Pooling: ÷2 Output/Dense 1: N=30; Act.=ReLU Output/Dense 2: N=3; Act.=Softmax
@@
1D-Justo-LucasCNN 25,323 Hidden: 1 1D Convolution & 1 Max Pooling Output: Flatten & 3 Dense Hidden/Convolution: N=16; K=9; Act.=Tanh Hidden/Pooling: ÷2 Output/Dense 1: N=30; Act.=Tanh Output/Dense 2: N=5; Act.=Tanh Output/Dense 3: N=3; Act.=Softmax
@@
2.3.3.2 Ablation Study
To adapt the soil models for deployment in space by reducing their trainable parameters while not degrading the models’ performance, we initially investigate the relevance of each neural layer and subsequently conduct hyperparameter fine-tuning. In the first column of Table I, we list the state-of-the-art soil models in italic font, while our customized models are listed in bold. We also differentiate our models by the addition of the term “Justo” to their names. In the subsequent columns of the table, for our models, we utilize bold font to highlight the hyperparameters that we modify compared to the soil models.
Taking into account the promising nature of 1D-CNNs as lightweight models for on-board deployment, we focus the hyperparameter optimization exclusively on them, and we omit the optimization for 2D-CNN models. The modifications to the hyperparameters in Table I further discussed in the following section, are a direct outcome of the hyperparameter fine-tuning process next. Namely, we begin by assigning them first intuitive values that can yield acceptable performance. Then, we conduct a search using a hyperparameter grid, based on the initial values, to identify the optimal combination of hyperparameters. To accommodate the high dimensionality of HSI data, we randomly choose only 2 million spectral signatures (approximately 10% of the training set) to reduce the time needed for the grid search. Additionally, while evaluating the grid models through cross-validation, we restrict the evaluation to only two fold permutations to decrease computation time further.
2.3.3.3 Introducing New Lightweigth 1D-CNNs:
@@
Figure 1: Lightweight 1D-Justo-LiuNet architecture for on-board segmentation. As an illustration, the block diagram employs 112 input channels for a single pixel. The network generates three detection probabilities, corresponding to sea, land, and clouds, respectively.
@@
The hyperparameter optimization process in previous section aims to achieve optimal accuracy, but we emphasize that the minimization of parameter counts for optimal in-orbit deployment is still the primary focus of our method. In this context, in our prior article [34], we presented a modified version of the Liuetal network resulting in 124,163 parameters, which might be considered large for in-flight deployment. However, our results demonstrated superior performance when contrasted with the Liuetal model [45]. We consider that the large parameter count in [34], which did not result in model overfitting, likely contributed to the significant performance improvement over the state-of-the-art. Nevertheless, in this article, our goal is to once again outperform the state-of-the-art, while also reducing the parameter count, in contrast to our previous work [34]. Indeed, we aim to achieve a parameter count even lower than the count of the Liuetal model in Table I. Therefore, we next present the 1D-Justo-LiuNet in Fig. 1, which achieves 80% reduction in parameters compared to Liuetal (22,755) [45]. 1D-Justo-LiuNet is adapted to only 4,563 parameters, as shown in Table I. This reduction is accomplished by first following the approach in our prior article [34], suggesting to start with a base seed of convolutional kernels and progressively increase the kernel counts (denoted as N in Table I) based on the seed multiplied by factors of 2, 3, and 4 as the network deepens. In this work, we however start with a small seed of 6 kernels, instead of 32 as in [34], to achieve the substantial 80% reduction in the model’s parameter count. Moreover, as we did in our previous publication, we double the kernel size (denoted as K in Table I) compared to [45] in order to compensate for the reduction in the kernel counts, thereby ensuring that performance is not adversely affected.
@@
For the Huetal model also in [45], the high parameter count in Table I is mainly a result of a fully connected layer with numerous neurons preceding the output classification layer. We refine the model, naming it 1D-Justo-HuNet, by reducing 70% the neural units in the fully connected layer. By implementing additional modifications such as reducing the number of convolutional kernels and adjusting their kernel size, we attain a significant 90.50% reduction in parameter count (9,543). As for LucasCNN [45], our model, 1D-Justo-LucasCNN, follows the same approach applied to Huetal to reduce the parameter count, i.e., we significantly reduce the neural units in the two dense layers before the classification layer. Specifically, LucasCNN has 280 neural dense units, while 1D-Justo-LucasCNN reduces the units by 87% to just 35 neurons. Furthermore, 1D-Justo-LucasCNN includes other adjustments, such as transforming the 4-scale LucasCNN into a single-scale network, significantly reducing the number of convolutional kernels while tripling the kernel size, and employing a hyperbolic neural activation instead of a linear function result of the hyperparameter optimization process. Table I illustrates a 68.41% reduction in parameter count (25,323) achieved by 1D-Justo-LucasCNN. The architectural details (as in Fig. 1) of the 1D-Justo-HuNet and 1D-Justo-LucasCNN are available in the supplementary materials, which also include the software codes to reproduce the models.
2.3.3.4 Introducing a New Lightweight 2D-CNN:
@@
Finally, to achieve lightweight models with reduced parameters, not only for 1D-CNN models as in previous section, but also for 2D-CNN architectures, we introduce: 2D-Justo-UNet-Simple with 7,641 parameters, as shown in Table I. The 2D-Justo-UNet-Simple consists of a compressed version of the encoder-decoder U-Net architecture with only 2 scales (as opposed to the typical 4 scales in U-Nets [40]), with a bottleneck level consisting of only 12 convolutional kernels. Additionally, following each 2D convolution operation and before pooling or upsampling, we insert batch normalization before the rectified linear activation to enhance model generalization and mitigate potential gradient vanishing or exploding effects during training. We note that in the decoder’s expanding path, we employ upsampling layers as an alternative to utilizing operations like transposed convolutions. While transposed convolutions may enhance performance, they also lead, however, to an increase in the number of trainable parameters. The software code and architectural details (as in Fig. 1) of the 2D-Justo-UNet-Simple are provided in the supplementary materials of this article.
2.3.4 Training Procedure
@@
We start by using the 112-channel dataset (from Section 2.2) to train the 1D and 2D models. However, when working with the 3-channel dataset (from Section 2.2 and A), we exclusively focus on training 2D models instead of 1D models. This decision is primarily based on the fact that 1D models do not take spatial information into account and rely solely on the spectral features, which, in this case, comprise only 3 channels. Furthermore, 1D-CNN models can face challenges related to the diminishing channel features as the network depth increases. Namely, when working with only three channels, the reduction in features is so pronounced that, in the first layers of the network, the processed features already diminish to zero, rendering prediction for three channels unfeasible.
In our training procedure, we initially select high-level parameters like the number of epochs and the batch size, followed by the choice of the lower-level hyperparameters previously discussed. In terms of high-level parameters, for both 1D-CNN and 2D-CNN models, we find that our dataset benefits from small batch sizes, increasing the frequency of updates in trainable parameters, but also extending the training time. As regards to batch size during inference, Keras dynamically selects the batch size for optimal memory efficiency. In addition, we find that it is advisable to utilize a small number of training epochs to prevent overfitting. Specifically, for 1D-CNNs, we use 2 epochs and a batch size of 32 pixels, while for 2D-CNNs, we opt for 3 epochs and a batch size of 4 patches. Furthermore, we employ the Adam optimizer with its default parameters to minimize the categorical cross-entropy loss, while using accuracy as evaluation metric in the training and validation sets.
@@
To evaluate the model’s performance at the end of each training epoch, we focus on the overall validation accuracy (validation set imbalance: 51.25%, 28.13%, and 20.62% for sea, land, and clouds, respectively), and we aim to achieve an empirical validation accuracy of around 0.90 during training. After completing the last training epoch, we evaluate false positives and false negatives (inversely proportional) by treating the segmentation of each class as a binary classification problem. In this context, we note that the significance of the false positives and false negatives varies per class, as elaborated further in the upcoming section. However, we note in advance that we aim for a reasonably balanced trade-off between false positives and false negatives across sea, land, and cloud classes before selecting the final trained model. Therefore, we consider false positives and false negatives intuitively during our training procedure, but we acknowledge more rigorous methods that focus on minimizing either false positives or false negatives by adjusting decision thresholds. The adjustments can involve, for instance, altering the decision boundary during the categorical class assignment based on different probability thresholds, consequently impacting the false positives and false negatives for each class. Additionally, applying penalties within the loss function to address specific class misclassifications can also affect false positives and false negatives. However, in our work, we consider sufficient to predict the categorical classes during inference by simply selecting the class with the highest probability, while equally penalizing all classes during training as the dataset imbalance is moderate. Finally, while quantization of trainable parameters certainly contributes to further compacting lightweight models [12], we emphasize that we do not perform parameter quantization in this article. The literature [15] demonstrates that in certain cases, where highly compressed light models are utilized (as in this work), quantization is unnecessary since the reduction in parameter count alone may be sufficient for in-orbit deployment. Furthermore, the effects of quantization can be highly dependent on the processing architecture.
@@
In the supplementary material of this article, we provide the trained model files with the learned parameters (46 distinct files). Additionally, we include a user-friendly software program we develop, offering an intuitive graphical user interface for conducting oceanic-terrestrial-cloud segmentation on HYPSO-1 imagery.
2.4 Segmentation Performance Metrics and

### v3

@@ II-C Semantic Segmentation Modeling
The models presented in this work handle the data in either one-dimensional (1D) or three-dimensional (3D) formats. On the one hand, when we refer to 1D processing, it implies that the models make predictions solely based on spectral characteristics, disregarding any nearby spatial context. Examples of such models include 1D-CNNs based on 1D convolutions, where the input consists of a spectral signature with no spatial context. Henceforth, we refer to these types of models as 1D models. However, when we refer to 3D processing, it signifies that the models concurrently analyze multiple adjacent spatial pixels, incorporating the hyperspectral datacube with both spectral (1D) and spatial (2D) information to predict the class for each pixel. Models such as 2D-CNNs, that we refer to henceforth as 2D models serve as examples, where the input is a 3D image containing both spatial and spectral data. Convolutional layers in 2D-CNN models operate by first applying 2D convolutions individually between each spectral channel and its respective neural kernel component. Subsequently, they aggregate the 2D convolutional results stacked in the spectral dimension to produce a final 2D feature map. Finally, we describe in Appendix B the arrangement of the data dimensions required for training and inference. In the appendix, we provide details on how the images are divided into non-overlapping 3D patches to train the 2D-CNNs and reduce computational complexity by processing each pixel only once, while using a small patch size of 48 $\times$ 48. This size ensures integer spatial dimensions through pooling layers halving dimensions of feature maps, while being also especially convenient for reducing computational complexity in orbit compared to processing larger patches.
II-C2 Baseline Segmentation Models
@@
As a starting point, we begin using simple baseline models consisting of traditional ML algorithms, as they are often more interpretable than deep neural networks. For convenience, during this initial phase, the models process 1D data (spectral signatures), which is a simpler processing strategy compared to handling 3D image patches. The initial ML baseline consists of Stochastic Gradient Descent (referred henceforth as 1D-ML/SGD), Gaussian Naïve Bayes (1D-ML/NB), Linear Discriminant Analysis (1D-ML/LDA), and Quadratic Discriminant Analysis (1D-ML/QDA). Next, we incorporate multiple state-of-the-art DL models for 3D patch processing, such as NU-Net [41] and NU-Net-mod [13], previously employed in satellite imagery for cloud detection in RGB (FACSAT-1) and MSI (Sentinel-2B). Furthermore, we incorporate additional DL models, namely C-UNet and C-UNet++, distinguished by the authors [42] as lightweight for their minimal number of parameters and low computational requirements, making them well-suited for deployment on satellites. C-UNet and C-UNet++ are compact versions of the standard encoder-decoder U-Net architecture [43]. The subsequent literature [44] building upon [42], explores C-UNet and C-UNet++ models suggesting their possible use for semantic segmentation in HS missions such as HYPSO-1. However, the method in [44] involves training on only two limited HS images, the conventional Pavia Centre and AeroRIT scene, which use sensors different from HYPSO-1 and do not include any cloud or oceanic data, which is of particular interest for ocean observation in HYPSO-1. In our work, we refer to the compact models in [44] as 2D-CUNet and 2D-CUNet++, and to their further reduced versions in [44] as 2D-CUNet Reduced and 2D-CUNet++ Reduced. Moreover, we add models from [45], which are referred to by Netteland [45] as the FAUBAI models, with reference to the FAUBAI project that focuses on segmentation for HS imagery for its possible future use aboard satellites. We refer to the networks in [45] as 2D-UNet FAUBAI and to its reduced version as 2D-UNet FAUBAI Reduced. Although Netteland [45] recommends any of these two networks for tasks such as cloud segmentation in HS satellite imagery, we note that similarly to [44], Netteland [45] conducted model testing only on a restricted set of four conventional HS images, namely, Pavia Centre, Pavia University, Salinas, and Indian Pines, which do not contain cloud or oceanic data.
Finally, we include one of the most recent and advanced lightweight models as a baseline. Fast Hybrid Vision Transformer (FastViT) is a state-of-the-art model introduced by Apple in 2023 [46] for processing image patches in tasks such as semantic segmentation, object detection, and image classification. FastViT is designed as a lightweight model optimized for edge inference on mobile devices and has been tested on platforms such as the iPhone 12 Pro. It introduces a novel token-mixing operator that utilizes structural reparameterization to enhance performance by eliminating skip connections, reducing memory access and footprint to optimize efficiency for mobile environments. Vasu et al. [46] proposed 8 architectures of various depths. From the set of architectures, we choose the shallowest one FastViT-T8, the deepest one (FastViT-MA36), and a middle-sized one (FastViT-S12).
II-C3 Customised Semantic Segmentation Models
II-C3.1   Review of 1D-CNNs: Liu, Hu, and Lucas CNNs
@@
The Liu [47], Hu [48] and Lucas [49] 1D-CNN models have been used before for data analysis in soil spectroscopy to segment various soil texture classes in HS data. Additionally, in our previous article [34], we verified that a modified version of LiuNet achieved high performance in satellite HS imagery, surpassing the state-of-the-art in soil spectroscopy [49]. In this work, we find LiuNet, HuNet, and LucasCNN of interest for in-orbit inference due to their simple 1D processing approach. Therefore, assuming 112 spectral channels and 3 classes to segment, in Table I, we show the parameter counts of these three models. The parameter counts are in the order of tens of thousands, with LiuNet having the lowest number of parameters among the three. Following our pursuit of further reducing the parameter count to merely a few thousands, more convenient for inference aboard satellites, in the next sections we accordingly modify the soil models in [49], while aiming to preserve their performance.
@@
LiuNet 22,755 Hidden: 4 1D Convolutions & 4 Max Poolings Output: Flatten & 1 Dense Hidden/Convolutions: N=32,32,64,64; K=3; Act.=ReLU Hidden/Poolings: ÷2 Output/Dense: N=3; Act.=Softmax
[…]
HuNet 100,663 Hidden: 1 1D Convolution & 1 Max Pooling Output: Flatten & 2 Dense Hidden/Convolution: N=20; K=12; Act.=Tanh Hidden/Pooling: ÷2 Output/Dense 1: N=100; Act.=Tanh Output/Dense 2: N=3; Act.=Softmax
[…]
@@
[…]
@@
II-C3.2   Ablation Study
To adapt the soil models for deployment in space by reducing their trainable parameters while not degrading their performance, we initially investigate the relevance of each neural layer and subsequently conduct hyperparameter fine-tuning. In the first column of Table I, we list the state-of-the-art soil models in italic font, while our customized models are listed in bold. We also differentiate our models by the addition of the term “Justo” to their names. In the subsequent columns of the table, for our models, we utilize bold font to highlight the hyperparameters that we modify with respect to the soil models.
Taking into account the promising nature of 1D-CNNs as lightweight models for on-board deployment, we focus the hyperparameter optimization exclusively on them, and we omit the optimization for 2D-CNN models. The modifications to the hyperparameters in Table I further discussed in the following section, are a direct outcome of the hyperparameter fine-tuning process. Namely, we begin by assigning them intuitive values that can yield acceptable performance. Then, we conduct a search using a hyperparameter grid, based on the initial values, to identify the optimal combination of hyperparameters. To accommodate the high dimensionality of HSI data, we randomly choose only 2 million spectral signatures (approximately 10% of the training set) to reduce the time needed for the grid search. Additionally, while evaluating the grid models through cross-validation, we restrict the process to only two fold permutations to further decrease computation time.
II-C3.3   Introducing New Lightweigth 1D-CNNs:
@@
The aforementioned hyperparameter optimization process aims to achieve optimal accuracy, but we emphasize that the minimization of parameter counts for optimal in-orbit deployment is still the primary focus of our method. In this context, in our prior article [34], we presented a modified version of the LiuNet network resulting in 124,163 parameters, which may be considered large for in-flight deployment. However, our results demonstrated superior performance when compared with the original LiuNet model [49]. We consider that the large parameter count in [34] likely contributed to the significant performance improvement over the state of the art. Nevertheless, in this article, our goal is to once again outperform the state of the art, while also reducing the parameter count, in contrast to our previous work [34].
Figure 1: Lightweight 1D-Justo-LiuNet architecture for on-board segmentation. The block diagram employs 112 input channels for a single pixel. The network generates three detection probabilities, corresponding to sea, land, and clouds, respectively.
Indeed, we aim to achieve a parameter count even lower than the count of the LiuNet model in Table I. Therefore, we next present the 1D-Justo-LiuNet in Fig. 1, with 80% reduction in parameters compared with LiuNet (22,755) [49]. 1D-Justo-LiuNet is reduced to only 4,563 parameters, as shown in Table I. This reduction is accomplished by first following the approach in our prior article [34], suggesting to start with a base seed of convolutional kernels and progressively increase the kernel counts (denoted as N in Table I) based on the seed multiplied by factors of 2, 3, and 4 as the network deepens. However, in this work, we start with a smaller seed of 6 kernels, instead of 32 as in [34], to achieve the substantial 80% reduction in the model’s parameter count. Moreover, as we did in our previous publication, we double the kernel size (denoted as K in Table I) compared with LiuNet [49] in order to achieve a larger receptive field for input samples and compensate for the reduced number of kernels, thereby ensuring that the performance is not adversely affected.
@@
For the HuNet model in [49], the high parameter count in Table I is mainly a result of a fully connected layer with numerous neurons preceding the output classification layer. We refine the model, naming it 1D-Justo-HuNet, by reducing 70% the neural units in the fully connected layer. By implementing additional modifications such as reducing the number of convolutional kernels and adjusting their kernel size, we attain a significant 90.50% total reduction in parameter count (9,543). Regarding LucasCNN [49], our model, 1D-Justo-LucasCNN, follows the same approach applied to HuNet to reduce the parameter count, i.e., we significantly reduce the neural units in the two dense layers before the classification layer. Specifically, LucasCNN has 280 neural dense units, while 1D-Justo-LucasCNN reduces the units by 87% to just 35 neurons. Furthermore, 1D-Justo-LucasCNN includes other adjustments, such as transforming the 4-scale LucasCNN into a single-scale network, significantly reducing the number of convolutional kernels while tripling the kernel size, and employing a hyperbolic neural activation instead of a linear function. Table I illustrates a 68.41% reduction in parameter count (25,323) achieved by 1D-Justo-LucasCNN. The architectural details (as in Fig. 1) of the 1D-Justo-HuNet and 1D-Justo-LucasCNN are available in our supplementary material, which also include the software codes to reproduce the models and an interactive Python notebook showing how model inference is performed.
II-C3.4   Introducing a New Lightweight 2D-CNN:
@@
Finally, to achieve lightweight models with reduced parameters, not only for 1D-CNN models as in previous section, but also for 2D-CNN architectures, we introduce 2D-Justo-UNet-Simple, a model with 7,641 parameters, as shown in Table I. The 2D-Justo-UNet-Simple consists of a compressed version of the encoder-decoder U-Net architecture with only 2 scales (as opposed to the standard 4-scale U-Net originally proposed for biomedical image segmentation in [43]), with a bottleneck level consisting of only 12 convolutional kernels. Additionally, following each 2D convolution operation and before pooling or upsampling, we insert batch normalization before the rectified linear activation to enhance model generalization and mitigate potential gradient vanishing or exploding effects during training. We note that in the decoder’s expanding path, we employ upsampling layers as an alternative to utilizing operations such as transposed convolutions. While transposed convolutions may enhance performance, they also lead, however, to an increase in the number of trainable parameters and complexity. The software code and architectural details (as in Fig. 1) of the 2D-Justo-UNet-Simple are also provided in the supplementary material of this article.
II-C4 Training Procedure
@@
We start by using the 112-channel dataset to train the 1D and 2D models. However, when working with the 3-channel dataset, we exclusively focus on training 2D models instead of 1D models. This decision is primarily based on the fact that 1D models do not take spatial information into account and rely solely on the spectral features, which, in this case, comprise only 3 channels. Furthermore, 1D-CNN models can face challenges related to the diminishing channel features as the network depth increases.
@@
In our training procedure, we initially select high-level parameters such as the number of epochs and the batch size, followed by the choice of the lower-level hyperparameters previously discussed. In terms of high-level parameters, for both 1D-CNN and 2D-CNN models, we find that our dataset benefits from small batch sizes, increasing the frequency of updates in trainable parameters, but also extending the training time. As regards to batch size during inference, Keras dynamically selects the batch size for optimal memory efficiency. In addition, we find that it is advisable to utilize a small number of training epochs to prevent overfitting. Specifically, for 1D-CNNs, we use 2 epochs and a batch size of 32 pixels, while for 2D-CNNs, we opt for 3 epochs and a batch size of 4 patches. Furthermore, we employ the Adam optimizer with its default parameters to minimize the categorical cross-entropy loss, while using accuracy as evaluation metric in the training and validation sets. To evaluate the model’s performance at the end of each training epoch, we focus on the overall validation accuracy (validation set imbalance: 51.25%, 28.13%, and 20.62% for sea, land, and clouds, respectively), and we aim to achieve an empirical validation accuracy of around 0.90 during training. After completing the last training epoch, we evaluate false positives and false negatives (inversely proportional) by treating the segmentation of each class as a binary classification problem. In our application, we note that the significance of the false positives and false negatives varies per class, as elaborated further in the upcoming section. However, we note in advance that we aim for a reasonably balanced trade-off between false positives and false negatives across sea, land, and cloud classes before selecting the final trained model. Therefore, we consider false positives and false negatives intuitively during our training procedure, but we acknowledge more rigorous methods that focus on minimizing either false positives or false negatives by adjusting decision thresholds. The adjustments can involve, for instance, altering the decision boundary during the categorical class assignment based on different probability thresholds, consequently impacting the false positives and false negatives for each class. Additionally, applying penalties within the loss function to address incorrect class predictions can also affect false positives and false negatives. However, in our work, we consider sufficient to predict the categorical classes during inference by simply selecting the class with the highest probability, while equally penalizing all classes during training as the dataset imbalance is moderate.
@@
In the supplementary material of this article, we provide the trained model files with the learned parameters (46 distinct files) for future in-orbit deployment. Additionally, we include a software data product with embedded inference by 1D-Justo-LiuNet within a user-friendly GUI for conducting oceanic-terrestrial-cloud segmentation on HYPSO-1 imagery.
II-D Segmentation Performance Metrics and